# Kreyòl-MT: Building MT for Latin American, Caribbean and Colonial African Creole Languages

### v1

@@ 1 Motivation
Public models achieving state-of-the-art performance on a published Creole language benchmark for 23 language directions
@@ 3.1 Collection methods
We divide our dataset collection methodology into two stages: gathering and extraction. Creole languages were selected for an active gathering effort (underlined in Table 1). We excluded Haitian because its data was abundant in already identified sources. Further data was found for other languages, also included in Table 1. For each of the gathering languages, we performed the following steps. First, we searched research databases using query templates to track down already curated datasets. We searched “[language name]” on the ACL Anthology,666https://aclanthology.org followed by “[language name] machine translation”, “[language name] NLP”, and “[language name] translation” on Google Scholar.777https://scholar.google.com Query results from these search engines were typically prohibitively many, so we browsed top results until it became clear that the remainder were no longer relevant. (Individual data gatherers used best judgment for each language.) For some languages we multiplied queries to accommodate for alternate language names. (See Table 8 in Appendix A) Second, for each language, we checked each of the following databases for parallel or monolingual corpora: OPUS (Tiedemann, 2012), Oscar (Ortiz Suárez et al., 2019) and LDC.888https://www.ldc.upenn.edu Third, we scoured the web for books with translations or additional resources. Fourth, we contacted researchers in the languages’ speaking communities for leads to potential data sources. Though we attempted various methods of contact whenever possible, response rates were low for this step. In all we gathered 107 sources such as anthology websites promoting cultural heritage or language revitalization, educational materials, and government documents.
@@
We immediately consolidated the previously published bitexts, including single-language resources from LAFAND-MT (Adelani et al., 2022a), KreolMorisienMT (Dabre and Sukhoo, 2022a) and the Caribe Trinidadian-English dataset (Smith, 2022). We gathered monolingual data from the MADLAD-400 clean corpus (Kudugunta et al., 2024) and JamPatoisNLI (Armstrong et al., 2022). To create novel bitexts, we then used the BeautifulSoup999https://www.crummy.com/software/BeautifulSoup/ and Selenium101010https://www.selenium.dev Python packages to extract text from web sources and the PyPDF2111111https://https://github.com/py-pdf/pypdf package for PDF sources. Organizing bitexts from sources with aligned parallel sentences was generally straightforward. When only unsegmented articles with translations were available, we segmented and aligned text based on punctuation and then manually corrected errors. Manual correction was performed by data extractors with sufficient proficiency in the languages involved. Our diverse team has some proficient speakers in our target languages (one L1 Martinican Creole speaker, two L2 Haitian speakers, one L2 Guadeloupean Creole speaker, one L2 Louisiana Creole speaker, two L2 Naija speakers, and one L2 Ghanaian Pidgin speaker). Though this did not cover anywhere near all of our included languages, proximity between Creole languages and related languages made the task doable.121212For instance, one Louisiana Creole speaker and one Haitian speaker—both also proficient in English, French, and linguistics—were able to manually correct alignment for Seychellois-English bitexts, due to Seychellois’ lexical overlap with Haitian, Louisiana Creole, and French. In general, the data extraction details varied for each source, since PDFs and websites have a wide variety of individualized styles and formats. In all, we extracted data from 39 of the 107 sources from our gathering stage.131313The remainder contained data in less accessible formats that we deemed too tedious to include in this work. We intend our dataset to be a living resource and hope to continually add to it as we encounter and format more data for Creole languages. Attributions for all data sources are in Appendix E.
@@ 4 Modeling Experiments
We now describe the experimental setup to show the utility of our datasets by training NMT systems. We train and release models on our full datasets, but we also train on only the data we publicly release, to show its stand-alone utility. We experiment with cleaning our train sets versus leaving them as-are, and in some experiments we fine-tune mBART (Liu et al., 2020) rather than training from scratch.
@@ 5 Results and Discussion
As expected, high-resource pairs exhibited better translation quality than their lower resource counterparts, with some deviations from this trend possibly due to inevitable non-uniformity in test set genre. Some pairs like Trinidadian-English exhibited chrF up to 84, and some zero-shot pairs like Annobonese-English scored as low as 0.
@@
Does fine-tuning help? From Figure 3, a comparison of scratch and mBART models trained on cleaned data (squares and x’s) reveals that, while models trained from scratch exhibit strong results, fine-tuning on mBART is better for both public and all data use. These fine-tuned models scored up to 8 chrF better than scratch models. Scratch models converged after 500K training steps of 33k-token batches, and mBART fine-tuning needed 500K-600K steps of only 8.2k-token batches. While this implies that mBART is data efficient, mBART fine-tuning itself is slow due to training about 700% more parameters. Indeed, when comparing the total training time, mBART fine-tuning needed up to a week whereas scratch training needed only up to two days on the same hardware.
Comparison on existing test sets: Figure 4 compares our best models against results reported by CreoleVal (Lent et al., 2023) on relevant language pairs from the CreoleVal test sets. Since CreoleVal data is not public, we consider it apt to evaluate our best model (mBART) fine-tuned on all cleaned data. Overall, our models surpass CreoleVal models with +6.4 average chrF XENG and +14.1 average chrF ENGX. Our dataset is much larger and more domain diverse than CreoleVal’s; hence our improved results on most language pairs show that increasing data tends to be beneficial even if it reduces the domain specificity. We note that on a few language directions (9 of 34), CreoleVal’s model still beat ours, indicating the possibility for negative interference from a more diverse train set in some instances. In Appendix C we provide some additional comparisons of bilingual versus multilingual fine-tuning and evaluations on other existing benchmarks: Lego-MT Yuan et al. (2023a) and KreolMorisienMT Dabre and Sukhoo (2022a).
@@ 6 Conclusion
In this work, we compile the most comprehensive dataset to date for MT of Creole languages in the Americas. By aggregating disparate previous works and incorporating new data sources via scraping the web and PDFs, we expand MT datasets to 21 new languages and produce the largest and most genre-diverse in 20 more. We release translation models in 172 language directions, with 23/32 beating state-of-the-art benchmark performance, as well as a public dataset with 11.6M aligned bitexts and 3.4M monolingual sentences.
@@ B.3 Hyperparameters
| data sampling temperature | 202020To slightly oversample the smaller corpora. We noticed in our peliminary experiments that the standard temperature of 5.0 caused the higher resource pairs to suffer. |
@@ B.4 Additional Models Trained
We train a simpler version of m2m-mBART focusing only on Haitian–English bidirectional translation to determine whether its better to focus on fewer language pairs during fine-tuning. The key changes in hyperparameters is learning rate () and dropout ().
@@ Appendix D Detailed Statistics on the Kreyòl-MT Dataset
| kmv | Karipuna | Amapá Creole | Uaçá Creole | - |
@@ E.1 Resources for Bitexts
E.2 Resources for Monolingual Corpora

### v2

@@ 1 Motivation
Public models achieving state-of-the-art performance on a published Creole language benchmark for 26 language directions
@@ 3.1 Collection methods
We divide our dataset collection methodology into two stages: gathering and extraction. Creole languages were selected for an active gathering effort (underlined in Table 1). We excluded Haitian because its data were abundant in already identified sources. Further data were found for other languages, also included in Table 1. For each of the gathering languages, we performed the following steps. First, we searched research databases using query templates to track down already curated datasets. We searched “[language name]” on the ACL Anthology,666https://aclanthology.org followed by “[language name] machine translation”, “[language name] NLP”, and “[language name] translation” on Google Scholar.777https://scholar.google.com Query results from these search engines were typically prohibitively many, so we browsed top results until it became clear that the remainder were no longer relevant. (Individual data gatherers used best judgment for each language.) For some languages we multiplied queries to accommodate for alternate language names. (See Table 8 in Appendix A.) Second, for each language, we checked each of the following databases for parallel or monolingual corpora: OPUS (Tiedemann, 2012), Oscar (Ortiz Suárez et al., 2019) and LDC.888https://www.ldc.upenn.edu Third, we scoured the web for books with translations or additional resources. Fourth, we contacted researchers in the languages’ speaking communities for leads to potential data sources. Though we attempted various methods of contact whenever possible, response rates were low for this step. In all we gathered 107 sources such as anthology websites promoting cultural heritage or language revitalization, educational materials, and government documents.
@@
We immediately consolidated the previously published bitexts, including single-language resources from LAFAND-MT (Adelani et al., 2022a), KreolMorisienMT (Dabre and Sukhoo, 2022a) and the Caribe Trinidadian-English dataset (Smith, 2022). We gathered monolingual data from the MADLAD-400 clean corpus (Kudugunta et al., 2024) and JamPatoisNLI (Armstrong et al., 2022). To create novel bitexts, we then used the BeautifulSoup999https://www.crummy.com/software/BeautifulSoup/ and Selenium101010https://www.selenium.dev Python packages to extract text from web sources and the PyPDF2111111https://https://github.com/py-pdf/pypdf package for PDF sources. Organizing bitexts from sources with aligned parallel sentences was generally straightforward. When only unsegmented articles with translations were available, we segmented and aligned text based on punctuation and then manually corrected errors. Manual correction was performed by data extractors with sufficient proficiency in the languages involved. Our diverse team has some proficient speakers in our target languages (one L1 Martinican Creole speaker, two L2 Haitian speakers, one L2 Guadeloupean Creole speaker, one L2 Louisiana Creole speaker, two L2 Naija speakers, and one L2 Ghanaian Pidgin speaker). Though this did not cover anywhere near all of our included languages, proximity between Creole languages and related languages made the task doable.121212For instance, one Louisiana Creole speaker and one Haitian speaker—both also proficient in English, French, and linguistics—were able to manually correct alignment for Seychellois-English bitexts, due to Seychellois’ lexical overlap with Haitian, Louisiana Creole, and French. In general, the data extraction details varied for each source, since PDFs and websites have a wide variety of individualized styles and formats. In all, we extracted data from 39 of the 107 sources from our gathering stage.131313The remainder contained data in less accessible formats that we deemed too tedious to include in this work. We intend our dataset to be a living resource and hope to continually add to it as we encounter and format more data for Creole languages. Attributions for all data sources are in Appendix D.
@@ 4 Modeling Experiments
We now describe the experimental setup to show the utility of our datasets by training MT systems. We train and release models on our full datasets, but we also train on only the data we publicly release, to show its stand-alone utility. We experiment with cleaning our train sets versus leaving them as-are, and in some experiments we fine-tune mBART (Liu et al., 2020) rather than training from scratch.
@@ 5 Results and Discussion
As expected, high-resource pairs exhibited better translation quality than their lower resource counterparts, with some deviations from this trend possibly due to inevitable non-uniformity in test set genre. Some possibly overfit pairs like Gullah-French scored chrF up to 97.2, and some zero-shot pairs like Papiamento-Spanish scored as low as 4.1.
@@
Does fine-tuning help? From Figure 3, a comparison of scratch and mBART models trained on cleaned data (squares and x’s) reveals that mBART fine-tuning helps both public and all data training. This fine-tuning helps more for out-of-Creole directions, with an average improvement of 8.1 chrF over scratch models on all data (compared to 1.4 chrF for into-Creole). Scratch models converged after 500k training steps of 33k-token batches, and mBART fine-tuning needed 500k-600k steps of only 8.2k-token batches. While this implies that mBART is data efficient, mBART fine-tuning is slow due to training about 700% more parameters. In total, mBART fine-tuning needed up to a week whereas scratch training needed only up to two days on the same hardware.
Comparison on existing test sets: Figure 4 compares our best models against results reported by CreoleVal (Lent et al., 2023) on relevant language pairs from the CreoleVal test sets. Since CreoleVal data is not public, we consider it apt to evaluate our best model (mBART) fine-tuned on all cleaned data. Overall, our models surpass CreoleVal models with +6.4 average chrF XENG and +14.1 average chrF ENGX. Our dataset is much larger and more domain diverse than CreoleVal’s; hence our improved results on most language pairs show that increasing data tends to be beneficial even if it reduces the domain specificity. We note that on a few language directions (8 of 34), CreoleVal’s model still beat ours, indicating the possibility for negative interference from a more diverse train set in some instances. In Appendix C we provide some additional comparisons of bilingual versus multilingual fine-tuning and evaluations on other existing benchmarks: Lego-MT Yuan et al. (2023a) and KreolMorisienMT Dabre and Sukhoo (2022a).
@@ 6 Conclusion
In this work, we compile the most comprehensive dataset to date for MT of Creole languages in the Americas. By aggregating disparate previous works and incorporating new data sources via scraping the web and PDFs, we expand MT datasets to 21 new languages and produce the largest and most genre-diverse in 20 more. We release translation models in 172 language directions, with 26/32 beating state-of-the-art benchmark performance, as well as a public dataset with 11.6M aligned bitexts and 3.4M monolingual sentences.
@@ B.3 Hyperparameters
| data sampling temperature202020We selected 2.0 to oversample the smaller corpora only slightly. We noticed in our peliminary experiments that the standard temperature of 5.0 caused the higher resource pairs to suffer. | |
@@ B.4 Additional Models Trained
We train a simpler version of m2m-mBART focusing only on Haitian–English bidirectional translation to determine whether its better to focus on fewer language pairs during fine-tuning. The key changes in hyperparameters are learning rate () and dropout ().
@@ C.3 Lego-MT and KreolMorisienMT Results
| kmv | Karipúna | Amapá Creole | Uaçá Creole | - |
@@ D.1 Resources for Bitexts
D.2 Resources for Monolingual Corpora